You shoot an arrow straight up at 50 m/s. When will it run out of speed?

Answer by stanbon(48557)

You shoot an arrow straight up at 50 m/s. When will it run out of speed?
------------
h(t) = -16y^2+50t
It will "run out of steam" when it's height is maximum.
That occurs when t = -b/2a = -50/[2(-16)] = 50/32 = 1.563 seconds
==========================
Cheers,
Stan H.